7 ( a ) Find The total resistances of 62 42 R,=? 52 121 512 62 (b) a 5.52 ba 112 Az < 4 Ras=? 5
Answer #1

br M 7 un Mb -Me 12r - M Mb 42 sr Ron + 1 / 2 = 6 R.6112 1276 6412 72 -402 18 12x6 R6,12 = 12+6 M hr Mh hor su Req = von Req=su 6 b M м а 6л. 52 bom ЛМ ИЛ а 3 Сл К%, 1| — 2л. Чл- Б Кіт, 11 = 6+ Г. 1 = 22л. 22x 2 so R4 21 +4 R6,22 bom — Ч. 21 л Ил Rab